Question 5
Whay the ans is not zero scince it is aclosed surface

the net electric flux is (a) 0, allu (U) dllu UU eo, A point charge of 1.8 is at the center of a cubical Gaussi surface 55 cm on edge. at is the net electric flux through the surface? Tssm
Answer #1

According to Gauss' Law,

you will get the same answer no matter where the charge q is placed inside the box, and no matter how large the box is. It does not even have to be a cube, as long as it encloses the charge.

Total flux = q/(epsilon-zero)

= 1.8*10^-6/8.82*10^-12

= 2.04*10^5 (N*m^2/C)

That is a lot easier than doing the actual integration of E(normal) over the six sides of the cube, but the answer will be the same.